jennylove:

20) Before a nuclear reaction, there is 0.0100 kg of uranium. After the nuclear reaction, there is 0.0090 kg of uranium. How much energy was released?

jennylove:

if someone can help me step by step ?

justjm:

I believe you must use the formula \[E=∆ m{c}^{2}\] and here \(∆m=0.0100kg-0.0090kg\) and \(c=2.998 \times 10^8 m s^{-1}\)

jennylove:

ohh okay, yes I remember that equation! thank you . that makes sense

justjm:

Yeah. You'll get units of \(kg~m^2~s^{-2}\) and that's the same as 1 J

jennylove:

wait it looks like the answer was :/ 4.50 × 10^13 J

jennylove:

(h = 6.63 × 10^-34 J • s) they provided me with this for the question as well if that helps any

jennylove:

i still want to figure out how to solve it to get the 4.50 *10^13 j , just for my understanding. do you have any idea?

justjm:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jennylove (h = 6.63 × 10^-34 J • s) they provided me with this for the question as well if that helps any \(\color{#0cbb34}{\text{End of Quote}}\) That's Planck's constant..so I'm assuming they solved this with Planck's formula \(E=hv\)

justjm:

v is frequency...is there anything about frequency provided? Or perhaps wavelength?

jennylove:

no, just the question that was given above^ , thats about it. hmm thats confusing :/

justjm:

Are you sure the answer is 4.50 × 10^13 J? I'm getting 8.928 * 10^13 J to 4 s.f.

jennylove:

thats what it says,unless theres an error:/ . how did you get 8.928 ?

jennylove:

oh wait no, im sorry yes its saying 9.0 × 10^13 J !

justjm:

\[E=∆mc^2\] \[E=(0.0100 kg - 0.0090 kg)(2.998 \times 10^8 m s^{-1})^2\] \[8.928 \times 10^{13}~kg~m^2~s^{-2}=8.928 \times10^{13} J\] to 4 s.f.

justjm:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jennylove oh wait no, im sorry yes its saying 9.0 × 10^13 J ! \(\color{#0cbb34}{\text{End of Quote}}\) lol it's fine xD

jennylove:

oh okay so you just multiply it be plancks constant

justjm:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jennylove oh okay so you just multiply it be plancks constant \(\color{#0cbb34}{\text{End of Quote}}\) Nah I didn't use Planck's constant, h, here. I'm using the constant for the speed of light. You would use Planck's constant in Planck's formula or some manipulation of it, which is \(E=hv\). The formula we used here to solve this problem is the well-known Einstein's formula

jennylove:

wait no , you did not use it. im sorry! i need to reread before I speak.

jennylove:

ohh okay , constant for speed! ill take note of that. let me reread everything and make sure i understand it. sorry about that

justjm:

All g

jennylove:

just solved it myself, and now im going to take note of it. thank you for your help . I appreciate it